# Experimental evidence for the physical delocalization of individual photons in an interferometer

### v1

@@ 1 Introduction
Recently, this traditional position has come under criticism, both on conceptual grounds and for practical reasons. Conceptually, Bell’s inequality violations have shown that quantum theory excludes certain models of reality, indicating that some statements about the past of a particle can be derived from the theoretical formalism [19, 20, 21, 22, 23, 24, 25]. On the practical side, new setups and new methods of measurement have provided experimental evidence for the non-classical aspects of single particle propagation in quantum mechanics [26, 27, 28, 31, 29, 30, 33, 34, 32]. A noteworthy breakthrough in this direction was the delayed-choice quantum eraser experiment, which showed that it is possible to “switch” between wave-like propagation (i.e. interference) and particle-like propagation (i.e. path information) by using photon entanglement[9, 10, 11, 12]. These results confirm that the wave-particle duality cannot be explained by a measurement independent reality. However, they fail to provide any evidence for the behaviors or individual photons. It is indeed difficult to obtain such information, since it must be extracted from the noisy statistics of many detection events. The trade-off relation between interference effects and paths means that any correlations between the two will be obscured by a considerable amount of background noise. Nevertheless it is possible to isolate such correlations in weak measurements, where the post-selection of a specific photon detection event allows us to evaluate conditional averages of complementary quantities [13, 14, 15]. The fact that these weak values lie outside the range of values observed in direct measurement has given rise to a long standing controversy regarding the physical meaning of weak measurements and their results [16, 18, 17, 35, 36, 37]. In a recent work, we therefore introduced a new method to quantify the statistical fluctuations of weak values by using quantum feedback to compensate the effect of a weak interaction between the system and a quantum probe [38]. This result was subsequently applied to demonstrate the delocalization of a single neutron between two paths of an interferometer, strongly indicating that weak values describe the delocalization of particles over the paths of an interferometer [39]. However, these results relied on an initial bias between the paths, making it difficult to apply them to the more conventional case of interferences between paths of equal amplitude. Here, we will address this problem and show that the delocalization of a single particle can be observed experimentally based on a direct observation of polarization flips as proposed in [40]. The experimental evidence reveals a non-trivial dependence of the observed delocalization on the precise phase shift, demonstrating the dependence of physical reality on the specific measurement applied to a quantum system.
@@ 2 Observable effects of photon delocalization
Figure 2: Method for observing photon delocalization. Vertically polarized photons are injected into a two-path interferometer. We apply local operations to the polarization using two Half-Wave-Plates(HWPs), HWP1 and HWP2, placed into the two paths of the interferometer. HWP1 rotates the polarization by a small angle $\theta_{0}(\ll 1)$, and HWP2 rotates the polarization in the opposite direction by the same angle $-\theta_{0}(\ll 1)$. Since the probability of a polarization flip from V-polarization to H-polarization is proportional to the square of the rotation angle, localized photons all flip with the same probability, $P(H|1)=P(H|2)$. When interference is observed, the polarization flip probabilities $P(H|\pm)$ observed in the output ports change, where lower flip probabilities indicate that the local rotations can cancel each other. The flip probabilities $P(H|\pm)$ thus provide direct evidence for the delocalization of photons inside the interferometer.
Here, we show that it is possible to obtain experimental evidence for this delocalization effect by evaluating the effects of local polarization rotations on the polarization observed in the output of the interferometer. For this purpose, we consider the interaction between the paths and the polarization quantitatively by assigning values of $+1$ to path 1 and $-1$ to path 2. The path quantity is then given by
@@
The quantity $\hat{A}$ represents the presence of the photon in the path 1 and path 2. If the initial state is the eigenstate |$\ket{1}$, $\hat{A}$ returns the eigenvalue $+1$, representing a photon localized in path 1. If the eigenstate is |$\ket{2}$, $\hat{A}$ gives the eigenvalue $-1$, representing a photon localized in path 2. The eigenvalues of $\hat{A}$ thus represent the localization of photons in the paths. Oppositely, any observation of values other than ±$\pm 1$ would be an indication that the photons are delocalized in some form.
We consider an input state represented by an equal superposition of the two paths, |$\ket{\psi}=\frac{1}{\sqrt{2}}(\ket{1}+\ket{2})$. If the photon is detected in one of the output ports of the interferometer, the photon cannot be detected in either of the paths and no eigenvalues can be assigned to the quantity $\hat{A}$. However, we can still obtain information about the value of $\hat{A}$ by transferring path information to the polarization of the photon. Starting with vertically polarized photons, we rotate the linear polarization by an angle of $\theta_{0}\hat{A}$, that is, we apply a local rotation of $+\theta_{0}$ in path 1 and a local rotation of $-\theta_{0}$ in path 2. Such polarization rotations can be realized by Half-Wave-Plates (HWPs) in the paths as indicated in Fig. 2. If the local angle $\theta_{0}$ is sufficiently small $(\ll 1)$, these rotations have very little effect on the interference observed in the output ports. The value of $\hat{A}$ is now encoded in the rotation angle of the photons, and the effects of this rotation by an angle of $\theta_{0}\hat{A}$ can be observed by detecting the polarization of the photons in the output of the interferometer.
@@
The physical effect of delocalization is a reduction of the polarization rotation and the associated flip probability $P(H|\pm)$. Oppositely, an increase of the flip probability $P(H|\pm)$ indicates an increase of the squared rotation angle $\theta^{2}$ beyond the local limit of $\theta_{0}^{2}$. This increase suggests that one of the two rotation angles must have changed its sign. This corresponds to a negative presence of the photon in one path, compensated by a presence greater than one in the other. The photon is “super-localized” in one of the paths, resulting in a difference between the contributions of the two paths that is greater than the one allowed for localized photons. In all cases, the magnitude of the quantity $\hat{A}$ is directly observed in the probability of the polarization flips, where values of $P(H|\pm)<\theta_{0}^{2}$ indicate delocalization and values of $P(H|\pm)>\theta_{0}^{2}$ indicate super-localization.
@@ 3 Experimental Setup
[…]
We found that the main source of experimental imperfections was the beam splitter through which the photons enter and exit the interferometer. The precise ratio of reflection and transmission at the beam splitter was 1.121, which corresponds to a reflectivity of 52.85 %. We note that this imbalance between transmission and reflection has a negligible effect on the visibility of the interference between the path. More problematic is the polarization dependence of this imbalance. To avoid errors in the measurement of delocalization, we blocked one path to characterize the rotated polarization of that path in the output. We could then confirm that the polarizations were rotated in opposite directions with a slight adjustment of the two GTs (details in the appendix A).
As explained in the previous section, photon polarization serves as a probe of the spatial delocalization of each photon between the paths. Specifically, the probability that a V-polarized input photon flips to H-polarization determines the value of $A^{2}(\pm)$ and therefore the delocalization of photons exiting the interferometer at the $+$ and the -- ports, respectively. Data was taken while varying the phase difference $\phi$ in 32 steps for a full range of $180\degree$, each step being about $5.625\degree$ in length. We switched the direction of the detected polarization between H and V, obtaining count rates for each polarization direction using the same time window 100s for each setting. This data was then used to evaluate the flip probabilities $P(H|+)$ and $P(H|-)$ when the photon is detected in a specific output port.
@@ 4 Results
We first characterized the interference fringes of the Sagnac-like interferometer by combining the counts of the H and V polarized photons at the output ports to obtain the total output probabilities $P(+)$ and $P(-)$. The interference fringes obtained in this measurement are shown in Fig. 4. Note that the polarization rotations in the paths have already been implemented. The experimentally observed visibility thus includes all decoherence effects caused by the interactions between polarizations and paths. The interference visibilities were evaluated to be 0.9575 for the $+$ output and 0.9629 for the -- output. The slight difference can be explained by the effects of imperfections of the beam splitter used in the interferometer, which have less impact on the $P(+)$ result at $\phi=0\degree$ than on the $P(-)$ result at $\phi=180\degree$.
@@
Figure 4: Probabilities $P(+)$ and $P(-)$ of detecting the photons in the respective output ports of the interferometer at different phases $\phi$. The phase was changed $-22.5\degree$ to $202.5\degree$ in steps of $5.625\degree$. The visibilities obtained from the data are 0.9575 for the $+$ output and 0.9629 for the -- output. These visibilities include the decoherence effects induced by the local polarization rotations in the paths.
@@
Fig. 5 and Fig. 6 show the experimental results for the flip probabilities $P(H|\pm)$ and the corresponding square of the path value $A^{2}(\pm)$. In Fig. 5, the polarization flip rates obtained for photons localized in one of the two paths are compared with the flip rates observed for photons observed after the two paths interfered. The data for the individual paths was taken by blocking the other path, while leaving all other settings unchanged. The results taken for photons localized in path 1 and in path 2 show that localized photons can only induce a polarization flip probability of about 0.015. The values obtained by averaging over the results obtained for different phase shifts are $0.01482\pm 0.00003$ for output port -- and $0.01587\pm 0.00003$ for output port $+$ (details in the appendix A). When both paths are open and interference occurs, the situation changes significantly, as shown by the solid circles in Fig. 5. As shown in Fig. 5 (a), a flip rate of $P(H|-)<0.015$ ($A^{2}(-)<1$) is observed for $\phi>90\degree$, indicating that the photons are physically delocalized between the paths whenever constructive interference favors the -- output. The flip probability $P(H|-)$ drops to nearly zero around $\phi=180\degree$, indicating that the polarization rotations in the two path cancel each other. This cancellation can only be achieved when each photon interacts equally with both HWP1 and HWP2. Quantitatively, the photons must be physically distributed over the two path, with half of the photon in one path, and half in the other. Likewise, as shown in Fig. 5 (b), a flip rate of $P(H|+)<0.015$ ($A^{2}(+)<1$) is observed for $\phi<90\degree$, indicating delocalization of the photons whenever constructive interference favors the $+$ output. The flip probability $P(H|+)$ drops to zero around $\phi=0\degree$, indicating maximal delocalization with half a photon in one path and half in the other.
@@
Figure 5: Delocalization of photons observed in output ports preferred by constructive interference. Graph (a) shows the phase dependence of $P(H|-)$, where constructive interference is observed for $\phi>90\degree$. Graph (b) shows the phase dependence of $P(H|+)$, where constructive interference is observed for $\phi<90\degree$. The solid circles represent the measurement results when the two paths interfered at the output. Solid squares and solid diamonds represent the data obtained when one of the paths was blocked. This data represents a value of $A^{2}(\pm)=1$, characteristic of localized photons. The axes on the right side of the graphs give the corresponding values of $A^{2}(\pm)$ based on this comparison. Delocalization is directly observed whenever constructive interference favors the output port in which the photon was detected.
@@
for all phases $\phi$. This relation has an interesting consequence for the upper limits of super-localization. Close to zero degrees, $A^{2}(+)$ is close to zero, but the probability $P(-)$ is also close to zero. It follows that, for $\phi=0$, $A^{2}(-)\approx 1/P(-)$. The maximal observable super-localization thus depends on the visibility of interference.
Figure 6: Super-localization of photons observed in output ports where the output probability is suppressed by destructive interference. Graph (a) shows the phase dependence of $P(H|-)$, where destructive interference is observed for $\phi<90\degree$. Graph (b) shows the phase dependence of $P(H|+)$, where destructive interference is observed for $\phi>90\degree$. The axes on the right side of the graphs give the corresponding values of $A^{2}(\pm)$. The flip probability $P(H|\pm)$ reaches its maximal value where the output probabilties $P(\pm)$ are minimal. $P(H|-)$ achieves a maximal value of $0.857\pm 0.005$ at $\phi=0\degree$ and $P(H|+)$ achieves a maximal value of $0.663\pm 0.002$ at $\phi=180\degree$. These maximal values are consistent with the visibilities of 0.9629 for the -- output and of 0.9575 for the $+$ output. The difference between the two maximum values can thus be explainable by the slight difference of the visibility of destructive interference in the two output ports. In terms of the $A^{2}(\pm)$, the maximal observed super-localizations are $A^{2}(-)=57.80\pm 0.34$ at $\phi=0\degree$ and $A^{2}(+)=41.78\pm 0.17$ at $\phi=180\degree$, roughly corresponding to a four-fold enhancement of the polarization rotation in one path, and a reversal and three-fold enhancement of the rotation in the other path.
Fig. 6 shows the experimental results for flip probabilities $P(H|\pm)>0.015$ ($A^{2}(\pm)>1$), indicating super-localization. As shown in Fig 6 (a), $P(H|-)$ peaks at a maximal value of $0.857\pm 0.005$ at $\phi=0\degree$, corresponding to a value of $A^{2}(-)=57.80\pm 0.34$. Eq.(3) associates this value of $A^{2}(-)$ with a probability of $P(-)=0.0173$, corresponding to a visibility of 0.965. The maximal value of $A^{2}(-)$ determined from the flip probabilities $P(H|-)$ is therefore consistent with the visibility of 0.9629 experimentally determined from the output probabilities $P(\pm)$ at $\phi=0\degree$. Fig 6 (b) shows the results for $P(H|+)$, where a maximal value of $0.663\pm 0.002$ is obtained at $\phi=180\degree$. The corresponding value of $A^{2}(+)$ is $41.78\pm 0.17$. The output probability estimated from the inverse $1/A^{2}(+)$ is 0.024, equivalent to a visibility of 0.952. The result is consistent with the visibility of 0.9575 experimentally determined from the output probabilities $P(\pm)$ at $\phi=180\degree$. The difference between the maximal values of super-localization at $\phi=0\degree$ and at $\phi=180\degree$ can thus be explained by the different visibilities observed at these phase shifts.

### v2

# arXiv:2505.00336v2[quant-ph] 27 Feb 2026

@@ 1 Introduction
Recently, this traditional position has come under criticism, both on conceptual grounds and for practical reasons. Conceptually, Bell’s inequality violations have shown that quantum theory excludes certain models of reality, indicating that some statements about the past of a particle could be derived from the theoretical formalism [9, 10, 11, 12, 13, 14, 15, 16, 17]. On the practical side, new setups and new methods of measurement have provided experimental evidence for the non-classical aspects of single particle propagation in quantum mechanics [18, 19, 20, 21, 22, 23, 24, 25, 26]. A noteworthy breakthrough in this direction was the delayed-choice quantum eraser experiment, which showed that it is possible to “switch” between wave-like propagation (i.e. interference) and particle-like propagation (i.e. path information) by using photon entanglement [27, 28, 29, 30]. These results confirm that the wave-particle duality cannot be explained by a measurement independent reality. However, they fail to provide any evidence for the behavior of individual photons. It is indeed difficult to obtain such information, since it must be extracted from the noisy statistics of many detection events. The trade-off relation between interference effects and paths means that any correlations between the two will be obscured by a considerable amount of background noise. Nevertheless it is possible to isolate such correlations in weak measurements, where the post-selection of a specific photon detection event allows us to evaluate conditional averages of complementary quantities [31, 32, 33, 34]. The fact that weak values lie outside the range of values observed in direct measurement has given rise to a long standing controversy regarding the physical meaning of weak measurements and their results [35, 36, 37]. A central problem is the understanding of the measurement process. In that regard, important insights have been gained through Hall’s discussion of Ozawa’s operator-based definition of uncertainties [38, 39]. Here, weak values appear as optimal estimates for each of the measurement results [40, 41, 42]. For pure states, Ozawa’s theory predicts a measurement uncertainty of zero when the measurement result is given by the weak value associated with a post-selected measurement result. Motivated by this observation, we recently introduced a method that directly quantifies the statistical fluctuations of weak values by using quantum feedback to compensate the effect of a weak interaction between the system and a quantum probe [43]. This result was subsequently applied to demonstrate the delocalization of a single neutron between two paths of an interferometer, strongly indicating that weak values describe the delocalization of particles over the paths of an interferometer [44]. However, these results relied on an initial bias between the paths, making it difficult to apply them to the more conventional case of interferences between paths of equal amplitude. Here, we will address this problem and show that the delocalization of a single particle can be observed experimentally based on a direct observation of polarization flips as proposed in [45]. The experimental evidence reveals a non-trivial dependence of the observed delocalization on the precise phase shift, demonstrating the dependence of physical reality on the specific measurement applied to a quantum system.
@@ 2 Observable effects of photon delocalization
Figure 2: Method for observing photon delocalization. Vertically polarized photons are injected into a two-path interferometer. We apply local operations to the polarization using two Half-Wave-Plates(HWPs), HWP1 and HWP2, placed into the two paths of the interferometer. HWP1 rotates the polarization by a small angle $\theta_{0}(\ll 1)$, and HWP2 rotates the polarization in the opposite direction by the same angle $\theta_{0}(\ll 1)$. Since the probability of a polarization flip from V-polarization to H-polarization is proportional to the square of the rotation angle, localized photons all flip with the same probability, $P(H|1)=P(H|2)$. When interference is observed, the polarization flip probabilities $P(H|\pm)$ observed in the output ports change, where lower flip probabilities indicate that the local rotations can cancel each other. The flip probabilities $P(H|\pm)$ thus provide direct evidence for the delocalization of photons inside the interferometer.
Here, we show that it is possible to obtain experimental evidence for this delocalization effect by evaluating the effects of local polarization rotations on the polarization observed in the output of the interferometer. For this purpose, we consider the interaction between the paths and the polarization quantitatively by assigning values of +1+1 to path 1 and 1-1 to path 2. The path quantity is then given by
@@
[…]
@@
The physical effect of delocalization is a reduction of the polarization rotation and the associated flip probability $P(H|\pm)$. Oppositely, an increase of the flip probability $P(H|\pm)$ indicates an increase of the squared rotation angle $\theta^{2}$ beyond the local limit of $\theta_{0}^{2}$. This increase suggests that one of the two rotation angles must have changed its sign. Since the interaction between path and polarization indicates that the polarization angle is proportional to the fraction of the particle present in the path, a change of rotation direction corresponds to a negative fraction of the photon in one of the paths. This negative fraction will be compensated by a corresponding fraction greater than one in the other path, resulting in an intensity difference between the paths that is greater than the photon number of one. When viewed as a single particle, the photon is effectively ”super-localized” in one of the paths, with a negative fractional presence in one path allowing for a much greater fluctuation in $\hat{A}$ than the limit for localized photons. In all cases, the magnitude of the quantity $\hat{A}$ is directly observed in the probability of the polarization flips, where values of $P(H|\pm)<\theta_{0}^{2}$ indicate delocalization and values of $P(H|\pm)>\theta_{0}^{2}$ indicate super-localization.
@@ 3 Experimental Setup
Our experimental setup is shown in Fig. 3. Its central part is a Sagnac-like interferometer that ensures the stability of the relative phase between the two paths. As photon source, we used a laser diode with a wavelength of $808.5\mbox{nm}$. The intensity was reduced by ND filters to a photon rate of 110000/s on average. The input photons enter and exit the interferometer through a 50:50 beam splitter (BS). Phase shifts between the paths were controlled using glass plates inserted into only one path each. The polarization of the output photons were discriminated with two Glan-Thompson Polarizers (GT) placed on each output just before a fiber coupler connected to an avalanche photo diode (APD) detector through an optical fiber. The dark counts of both APDs were evaluated by blocking the light from the photon source, allowing us to subtract the dark count background from the total counts to achieve accurate count rates at low detection probabilities (details in Appendix B).
We found that the main source of experimental imperfections was the beam splitter through which the photons enter and exit the interferometer. The precise ratio of reflection and transmission at the beam splitter was 1.121, which corresponds to a reflectivity of 52.85 %. We note that this imbalance between transmission and reflection has a negligible effect on the visibility of the interference between the path. More problematic is the polarization dependence of this imbalance. To avoid errors in the measurement of delocalization, we blocked one path to characterize the rotated polarization of that path in the output. We could then confirm that the polarizations were rotated in opposite directions with a slight adjustment of the two GTs (details in Appendix A1.2).
As explained in the previous section, photon polarization serves as a probe of the spatial delocalization of each photon between the paths. Specifically, the probability that a V-polarized input photon flips to H-polarization determines the value of $A^{2}(\pm)$ and therefore the delocalization of photons exiting the interferometer at the ++ and the - ports, respectively. Data was taken while varying the phase difference $\phi$ in 32 steps for a full range of $180\degree$, each step being about $5.625\degree$ in length. We switched the direction of the detected polarization between H and V, obtaining count rates for each polarization direction using the same time window 100s for each setting. This data was then used to evaluate the flip probabilities $P(H|+)$ and P(H|)P(H|-) when the photon is detected in a specific output port.
@@ 4 Results
We first characterized the interference fringes of the Sagnac-like interferometer by combining the counts of the H and V polarized photons at the output ports to obtain the total output probabilities $P(+)$ and P()P(-). The interference fringes obtained in this measurement are shown in Fig. 4. Note that the polarization rotations in the paths have already been implemented. The experimentally observed visibility thus includes all decoherence effects caused by the interactions between polarizations and paths. The interference visibilities were evaluated to be 0.9575 for the ++ output and 0.9629 for the - output. The slight difference can be explained by the effects of imperfections of the beam splitter used in the interferometer, which have less impact on the $P(+)$ result at $\phi=0\degree$ than on the P()P(-) result at $\phi=180\degree$.
@@
Figure 4: Probabilities $P(+)$ and P()P(-) of detecting the photons in the respective output ports of the interferometer at different phases $\phi$. The phase was changed $22.5\degree$ to $202.5\degree$ in steps of $5.625\degree$. The visibilities obtained from the data are 0.9575 for the ++ output and 0.9629 for the - output. These visibilities include the decoherence effects induced by the local polarization rotations in the paths.
@@
Fig. 5 and Fig. 6 show the experimental results for the flip probabilities $P(H|\pm)$ and the corresponding square of the path value $A^{2}(\pm)$. In Fig. 5, the polarization flip rates obtained for photons localized in one of the two paths are compared with the flip rates observed for photons observed after the two paths interfered. The data for the individual paths was taken by blocking the other path, while leaving all other settings unchanged. The results taken for photons localized in path 1 and in path 2 show that localized photons can only induce a polarization flip probability of about 0.015. The values obtained by averaging over the results obtained for different phase shifts are $0.01482\pm 0.00003$ for output port - and $0.01587\pm 0.00003$ for output port ++ (details in Appendix A1.2). When both paths are open and interference occurs, the situation changes significantly, as shown by the solid circles in Fig. 5. As shown in Fig. 5 (a), a flip rate of P(H|)<0.015P(H|-)<0.015 ($A^{2}(-)<1$) is observed for $\phi>90\degree$, indicating that the photons are physically delocalized between the paths whenever constructive interference favors the - output. The flip probability P(H|)P(H|-) drops to nearly zero around $\phi=180\degree$, indicating that the polarization rotations in the two path cancel each other. This cancellation can only be achieved when each photon interacts equally with both HWP1 and HWP2. Quantitatively, the photons must be physically distributed over the two path, with half of the photon in one path, and half in the other. Likewise, as shown in Fig. 5 (b), a flip rate of $P(H|+)<0.015$ ($A^{2}(+)<1$) is observed for $\phi<90\degree$, indicating delocalization of the photons whenever constructive interference favors the ++ output. The flip probability $P(H|+)$ drops to zero around $\phi=0\degree$, indicating maximal delocalization with half a photon in one path and half in the other.
@@
Figure 5: Delocalization of photons observed in output ports preferred by constructive interference. Graph (a) shows the phase dependence of P(H|)P(H|-), where constructive interference is observed for $\phi>90\degree$. Graph (b) shows the phase dependence of $P(H|+)$, where constructive interference is observed for $\phi<90\degree$. The solid circles represent the measurement results when the two paths interfered at the output. Solid squares and solid diamonds represent the data obtained when one of the paths was blocked. This data represents a value of $A^{2}(\pm)=1$, characteristic of localized photons. The axes on the right side of the graphs give the corresponding values of $A^{2}(\pm)$ based on this comparison. Delocalization is directly observed whenever constructive interference favors the output port in which the photon was detected.
@@
for all phases $\phi$. This relation has an interesting consequence for the upper limits of super-localization. Close to zero degrees, $A^{2}(+)$ is close to zero, but the probability P()P(-) is also close to zero. It follows that, for $\phi=0$, $A^{2}(-)\approx 1/P(-)$. The maximal observable super-localization thus depends on the visibility of interference.
Figure 6: Super-localization of photons observed in output ports where the output probability is suppressed by destructive interference. Graph (a) shows the phase dependence of P(H|)P(H|-), where destructive interference is observed for $\phi<90\degree$. Graph (b) shows the phase dependence of $P(H|+)$, where destructive interference is observed for $\phi>90\degree$. The axes on the right side of the graphs give the corresponding values of $A^{2}(\pm)$. The flip probability $P(H|\pm)$ reaches its maximal value where the output probabilties $P(\pm)$ are minimal. P(H|)P(H|-) achieves a maximal value of $0.857\pm 0.005$ at $\phi=0\degree$ and $P(H|+)$ achieves a maximal value of $0.663\pm 0.002$ at $\phi=180\degree$. These maximal values are consistent with the visibilities of 0.9629 for the - output and of 0.9575 for the ++ output. The difference between the two maximum values can thus be explainable by the slight difference of the visibility of destructive interference in the two output ports. In terms of the $A^{2}(\pm)$, the maximal observed super-localizations are $A^{2}(-)=57.80\pm 0.34$ at $\phi=0\degree$ and $A^{2}(+)=41.78\pm 0.17$ at $\phi=180\degree$, roughly corresponding to a four-fold enhancement of the polarization rotation in one path, and a reversal and three-fold enhancement of the rotation in the other path.
Fig. 6 shows the experimental results for flip probabilities $P(H|\pm)>0.015$ ($A^{2}(\pm)>1$), indicating super-localization. As shown in Fig. 6 (a), P(H|)P(H|-) peaks at a maximal value of $0.857\pm 0.005$ at $\phi=0\degree$, corresponding to a value of $A^{2}(-)=57.80\pm 0.34$. Eq.(3) associates this value of $A^{2}(-)$ with a probability of P()=0.0173P(-)=0.0173, corresponding to a visibility of 0.965. The maximal value of $A^{2}(-)$ determined from the flip probabilities P(H|)P(H|-) is therefore consistent with the visibility of 0.9629 experimentally determined from the output probabilities $P(\pm)$ at $\phi=0\degree$. Fig. 6 (b) shows the results for $P(H|+)$, where a maximal value of $0.663\pm 0.002$ is obtained at $\phi=180\degree$. The corresponding value of $A^{2}(+)$ is $41.78\pm 0.17$. The output probability estimated from the inverse $1/A^{2}(+)$ is 0.024, equivalent to a visibility of 0.952. The result is consistent with the visibility of 0.9575 experimentally determined from the output probabilities $P(\pm)$ at $\phi=180\degree$. The difference between the maximal values of super-localization at $\phi=0\degree$ and at $\phi=180\degree$ can thus be explained by the different visibilities observed at these phase shifts.